OpenStudy (anonymous):

An average person in the United States throws away 4.5 × 10–1 kg of trash per day. A single gram of trash has an average volume of 3 cm3. How many cubic meters of trash would the average person in the United States produce each year (1 m3 = 1 × 106 cm3)?

OpenStudy (anonymous):

298.08 meters per hour

OpenStudy (anonymous):

A. 2.76 × 108 m/h B. 2.76 × 1010 m/h C. 7.67 × 106 m/h D. 7.67 × 108 m/h

OpenStudy (anonymous):

please answer! Anyone!

OpenStudy (anonymous):

will medal and fan

OpenStudy (tkhunny):

1) Why are the answers written like that? Very, VERY confusing. Perhaps it should be \(4.6 * 10^{8}\;cm/min\)? 2) Didn't we already do this? 2a) cm ==> m multiply the expression by \(\dfrac{1\;m}{100\;cm}\) 2b) min ==> hr multiply the expression by \(\dfrac{60\;min}{1\;hr}\) Let's see what you get.

OpenStudy (anonymous):

29.08 meters per hour but i need in scientific notation expression

OpenStudy (anonymous):

An average person in the United States throws away 4.5 × 10–1 kg of trash per day. A single gram of trash has an average volume of 3 cm3. How many cubic meters of trash would the average person in the United States produce each year (1 m3 = 1 × 106 cm3)?

OpenStudy (tkhunny):

Not sure where you managed 29.08 4.6*60/100 = 2.76 The magnitude is still \(10^{8}\)

OpenStudy (anonymous):

i still dont get the answer

OpenStudy (tkhunny):

You may wish to stick with one problem at a time. Try a) 2.76 * 10^8

OpenStudy (anonymous):

An average person in the United States throws away 4.5 × 10–1 kg of trash per day. A single gram of trash has an average volume of 3 cm3. How many cubic meters of trash would the average person in the United States produce each year (1 m3 = 1 × 106 cm3)?

OpenStudy (anonymous):

is anyone gonna help?

OpenStudy (tkhunny):

It works better if you post new problems on new threads. Did you show any of your OWN work? That helps, too.